# Fitts’ List Revisited: An Empirical Study on Function Allocation in a Two-Agent Physical Human-Robot Collaborative Position/Force Task

### v1

@@ 4 Discussion
This study investigated the relevance of Fitts’ classical “Men Are Better At / Machines Are Better At” (MABA-MABA) framework [11] in the context of two-agent pHRC. The central question aimed to identify the optimal static allocation of position and force control between a human and a robot. The findings presented here provide compelling evidence suggesting that the fundamental principles articulated by Fitts over seven decades ago retain significant relevance for designing effective collaboration strategies in modern pHRC scenarios involving direct physical interaction and shared control. Beyond this confirmation, the results also yield additional important insights into user experience.
4.1 Validating Fitts: HR outperforms RH
@@
The direct comparison between the HR (Human Position, Robot Force) and RH (Robot Position, Human Force) conditions consistently favored HR across multiple dimensions. Participants reported higher acceptance (supporting H1, Fig. 3), and demonstrated significantly less overblending, while completion time was found to be similar (largely supporting H2, Fig. 5). They also experienced significantly less frustration and lower physical demand (largely supporting H3, Fig. 4), and experienced greater autonomy, felt more useful, and engaged (strongly supporting H4, Fig. 6) when they controlled positioning while the robot controlled the force. These findings align well with the “MABA-MABA” principle: leveraging human dexterity and adaptability for position control and delegating force exertion to the machine leads to a more effective and preferred collaboration.
4.2 RH leads to increased overblending and frustration
@@
The difficulties encountered in the RH condition, notably increased overblending and elevated frustration, suggest a misalignment. This could stem from the robot dictating the position trajectory, potentially introducing unpredictability, or a path that conflicts with the human’s natural movement tendencies or intentions for force application. This most likely forces the human operator to be reactive, attempting to synchronize their force output with a robot-imposed motion profile. A dip in self-rated competence (although not found to be statistically significant, see Fig. 6) could result from a loss in confidence due to this misalignment.
4.3 HR provides better experience compared to HH
@@
Interestingly, the HR condition was not only preferred over RH, but was also rated significantly more useful and satisfying than the HH baseline (Fig. 3). This suggests that the Fitts-aligned collaboration is not just relatively better but that HR can provide a genuinely enhanced experience compared to unassisted manual work. This is likely due to the substantial reduction in physical demand (Fig. 4), without compromising the user’s sense of agency over the task’s positional execution or a large drop in perceived autonomy (see HH vs HR and RH in Fig. 6).
@@
It is important to note that neither HR nor RH improved objective task performance compared to HH (Fig. 5). On a surface level, this might hint that robots are not necessary since there is no performance increase with a robot. Nevertheless, pHRC has many other benefits that do not directly affect the performance. For example, pHRC improves ergonomics in physically demanding tasks, such as blending, to prevent work-related musculoskeletal injuries [16].
4.4 Supervisory control and the ironies of automation
The RR condition presented a unique profile characterized by trade-offs. Objectively, RR showed the best completion times and the lowest cumulative overblended health, as it was designed to do (Fig. 5). It also resulted in the lowest perceived workload across almost all NASA-TLX dimensions (Fig. 4). However, these efficiency gains came at a cost: participants reported the lowest levels of perceived autonomy, usefulness, and engagement (Fig. 6). This resonates with Bainbridge’s classic “ironies of automation” [25]: optimizing system performance can paradoxically diminish human experience by reducing meaningful involvement.
Nevertheless, our findings, specific to pHRC, also partially contradict that classical notion since RR achieved surprisingly favorable acceptance ratings, ranking second overall, statistically comparable to HH, though still lower than HR (Fig. 3). One explanation is that the substantial reduction in physical and temporal demands was highly valued by participants. Moreover, the supervisory role, requiring vigilance and occasional intervention, likely provided more engagement (perceived similarly to RH, see Fig. 6) than purely passive oversight, mitigating the worst effects of automation. Participants also rated the robot in RR as more useful and engaging than in HH and RH (Fig. 6), suggesting they recognized its effectiveness even while acknowledging their own reduced role.
@@ 4.4 Supervisory control and the ironies of automation
4.5 The primacy of position control for experienced autonomy
Results reveal a pronounced asymmetry in how delegation affected participants’ sense of autonomy (Fig. 6). Delegating position control to the robot (HH vs RH) resulted in a substantially larger decrease in reported autonomy compared to relinquishing force control (HH vs HR). Suggesting that position control is more intrinsically linked to the operator’s sense of agency than force control.
Position control determines where the interaction occurs and directs the tool’s action towards achieving the task goal, whereas force control modulates the intensity of the interaction at a given location. Participants may perceive the goal-directed aspect as the primary locus of control in executing the physical task. Relinquishing positional authority may thus feel akin to losing command over task progression, downgrading the human contribution to one of reactive effort modulation–analogous to adjusting the throttle as a passenger, rather than steering the vehicle as a driver.
This autonomy drop aligns with Self-Determination Theory (SDT) [22], which identifies autonomy as a psychological need that supports motivation and engagement. Work designs that constrain agency, are likely to thwart this need and reduce self-determined motivation. Indeed, the RH condition not only lowered perceived autonomy but also significantly decreased perceived engagement and usefulness. These findings support the SDT proposition that motivation and psychological well-being depend on environments that support, rather than undermine, autonomy.
4.6 Introduction of robot does not affect mental demand
The lack of significant difference in mental demand across conditions (Fig. 4) is intriguing. It might suggest that each condition imposed a comparable cognitive load, albeit of a different nature: HH required focused motor control for both sub-tasks, HR required coordination with robot force, RH demanded a more reactive force modulation, and RR necessitated supervisory vigilance. These results suggest that introducing a robot in any function does not make the task more mentally demanding for humans, which can be a valuable insight for workplaces interested in pHRC solutions.

### v2

@@ IV Discussion
This study investigated the relevance of Fitts’ classical “Men Are Better At / Machines Are Better At” (MABA-MABA) framework [fitts_human_1951] in the context of two-agent pHRC. The central question aimed to identify the optimal static allocation of position and force control between a human and a robot during physical collaboration. The findings presented here provide compelling evidence suggesting that the fundamental principles articulated by Fitts over seven decades ago retain significant relevance for designing effective collaboration strategies in modern pHRC scenarios involving direct physical interaction and shared control. Beyond this validation, the results also yield additional insights into user experience.
@@
Validating Fitts. HR outperforms RH: The direct comparison between the HR and RH conditions consistently favored HR across multiple dimensions. Participants reported higher acceptance (supporting H1, Fig. 3), and demonstrated significantly less overblending, while completion time was found to be similar (largely supporting H2, Fig. 5). This equivalence in completion time likely reflects a balance of two different inefficiencies: the time lost to the human’s imperfect exploratory path in HR was matched by the time the robot in RH spent on corrective re-blending due to the human’s imperfect force application for this specific task. Participants also experienced significantly less frustration and lower physical demand (largely supporting H3, Fig. 4), and experienced greater autonomy, felt more useful, and engaged (strongly supporting H4, Fig. 6) when they controlled positioning while the robot controlled the force. These findings align well with the “MABA-MABA” principle: leveraging human dexterity and adaptability for position control and delegating force to the machine leads to a more effective and preferred collaboration.
@@
RH leads to increased overblending and frustration: The difficulties encountered in the RH condition, notably increased overblending and elevated frustration, suggest a misalignment. This could stem from the robot dictating the position trajectory, potentially introducing unpredictability, or a path that conflicts with the human’s natural movement tendencies or intentions. This most likely forces the human operator to be reactive, attempting to synchronize their force output with a robot-imposed motion profile. A dip in self-rated competence (although not found to be statistically significant, see Fig. 6) could result from a loss in confidence due to this misalignment.
@@
HR provides better experience compared to HH: Interestingly, the HR condition was not only preferred over the inverse task allocation (RH), but was also rated as significantly more useful and satisfying than the HH baseline (Fig. 3), suggesting the Fitts-aligned collaboration (HR) provides a genuine enhancement over unassisted work. This is likely due to the substantial reduction in physical demand (Fig. 4), without compromising the user’s sense of agency over the task’s positional execution or a large drop in perceived autonomy (see HH vs HR and RH in Fig. 6).
@@
It is important to note that neither HR nor RH improved objective task performance compared to HH (Fig. 5). The absence of a performance increase could imply that robot assistance is superfluous. Nevertheless, pHRC has many other benefits that do not directly affect the performance. For example, pHRC improves ergonomics in physically demanding tasks, such as blending, to prevent work-related musculoskeletal injuries [peternel_robot_2018].
Supervisory control and the ironies of automation: The RR condition presented a unique profile characterized by trade-offs. Objectively, RR showed the best completion times and the lowest cumulative overblended health, as it was designed to do (Fig. 5). It also resulted in the lowest perceived workload across almost all NASA-TLX dimensions (Fig. 4). However, these efficiency gains came at a cost: participants reported the lowest levels of perceived autonomy, usefulness, and engagement (Fig. 6). This resonates with Bainbridge’s classic “ironies of automation” [bainbridge_ironies_1983]: optimizing system performance can paradoxically diminish human experience by reducing meaningful involvement.
Nevertheless, our findings, specific to pHRC, also partially contradict that classical notion since RR achieved surprisingly favorable acceptance ratings, ranking second overall, statistically comparable to HH, though still lower than HR (Fig. 3). One explanation is the substantial reduction in physical and temporal demands was highly valued by participants. Moreover, the supervisory role, requiring vigilance and occasional intervention, likely provided more engagement (perceived similarly to RH, see Fig. 6) than purely passive oversight, mitigating the worst effects of automation. Participants also rated the robot in RR as more useful and engaging than in HH and RH (Fig. 6), suggesting they recognized its effectiveness even while acknowledging their own reduced role.
@@
The primacy of position control for experienced autonomy: Results reveal a pronounced asymmetry in how delegation affected participants’ sense of autonomy (Fig. 6). Delegating position control to the robot (HH vs RH) resulted in a substantially larger decrease in reported autonomy compared to relinquishing force control (HH vs HR), suggesting that position control is more intrinsically linked to the operator’s sense of agency than force control.
Position control determines where the interaction occurs and directs the tool’s action towards achieving the task goal, whereas force control modulates the intensity of the interaction at a given location. Participants may perceive the goal-directed aspect as the primary locus of control in executing the physical task. Relinquishing positional authority may thus feel akin to losing command over task progression, downgrading the human contribution to one of reactive effort modulation.
This drop in autonomy when delegating position control aligns well with Self-Determination Theory (SDT) in a pHRC context. SDT posits that autonomy is a fundamental psychological need, and its frustration thwarts motivation and well-being, which can lead to negative outcomes such as burnout. Our findings empirically mirror this principle: RH recorded the lowest levels of perceived autonomy (shared with RR) and also yielded the highest levels of frustration and the lowest satisfaction scores. This suggests that for participants, delegating position control was not just a change in task mechanics but the removal of a psychological need, which resulted in a negative experience.
Introduction of robot does not affect mental demand: The non-significant difference in mental demand suggests a comparable cognitive load, albeit of a different nature: HH required focused motor control for both sub-tasks, HR required coordination with the robot producing force, RH demanded a more reactive vigilance to synchronize with the robot’s movement, and RR necessitated supervisory vigilance. This suggests that introducing a robot in any function does not make the task largely more mentally demanding, which can be a valuable insight for workplaces considering pHRC solutions.
Limitations: Several limitations warrant consideration. The predominantly male participant pool (4 female, 22 male) restricts generalisability, and the abstracted lab task may not fully transfer to complex industrial settings where optimal task allocation could differ. Furthermore, our evaluation was based on short-term interaction; long-term use, especially in the supervisory (RR) condition, could alter user experience due to factors like sustained vigilance. Finally, as the new ACES questionnaire lacks full psychometric validation, its findings should be interpreted with caution.
Acknowledgment
The authors would like to thank Alexis Derumigny for advice on the statistical evaluation.